A regular pentagon can be dissected into 5 isosceles triangles whose vertex angle is at the center of the pentagon. The height of the triangle is 10 cm. find the area of this pentagon.: A regular pentagon can be dissected into 5 isosceles triangles whose vertex angle is at the center of the pentagon. The height of the triangle is 10 cm. find the area of this pentagon.
Answer by mangopeeler07(445)

Pentagon=540 degrees
540/5=108
108=the vertex angle of the isosceles triangle

1/2 of 108=54
tan(54)=x/10
x=13.764
(x=length of half of one side of the pentagon)
2x=27.528

apothem=10
perimeter=137.638

area=1/2ap

137.638(10)=1376.38

1376.38/2=688.19

area=688.19cm^2